Consider an automobile moving at v mph that skids d feet after its brakes lock. Calculate how far it would skid if it was moving at 2v and the brakes were locked.








1
Expert's answer
2021-12-06T09:45:54-0500

Assuming the same acceleration in both cases, first, find the deceleration:


"a=v^2\/(2d_v)."


Now, find the distance upon the same acceleration but a greater speed:


"d_{2v}=(2v)^2\/(2a)=4v^2\/(v^2\/d)=4d_v."